Subject: APO 3.5m users committee minutes, April 2012

From: Suzanne Hawley

Submitted: Mon, 16 Apr 2012 23:18:43 -0700 (PDT)

Message number: 1269 (previous: 1268, next: 1270 up: Index)

 		APO 3.5-m Users Committee Phonecon, 4/9/12

Attending:  Suzanne Hawley, Mark Klaene, Bill Ketzeback, Sean Moran, 
Jon Holtzman, John Bally, Michael Strauss, Al Harper, Remy Indebetouw, 
Scott Anderson

Absent: Bruce Gillespie


User feedback and comments from institutional representatives:

   o Chicago (Al Harper) - nothing to report.

   o Washington (Scott Anderson) - nothing to report.

   o New Mexico State (Jon Holtzman) - nothing to report.

   o Princeton (Michael Strauss) - Recent issues with triplespec guiding,
probably due to mask problems.

   o Colorado (John Bally) - Question about Q2 passwords, APO staff will

   o Virginia (Remy Indebetouw) - nothing to report.

   o Johns Hopkins (Sean Moran) - User had trouble downloading data
from JHU Applied Physics Lab computer


o telescope/instruments report: Mark presented the following report,
there were no questions.


   	      3.5-m Telescope and Instruments Highlights
   	                 2/23/12 through 04/04/12

0) Overview

Telescope continued to run well, and we had some great weather.  Some 
minor instrument related problems occurred. Two APO staff attended a 
training seminar on maintenance and operations of the Polycold Cryotigers. 
With a lot of effort by staff and UWy team and some help from Matt Nelson 
at UVa the NIIS run was successful in acquiring some first light images.

1) Telescope

The telescope has been operational with no issues of note. Tracking issues 
showed up for a partial night. They seem to have corrected themselves but 
are being investigated.

2) Instruments

DIS replacement Blue 400 grating was installed and aligned and is now 
available for routine observations.  Difference between new and old 
grating is negligible.  We did have another instance of vacuum level 
rising; a quick pump down allowed the instrument to continue observing 
until a new ION pump could be installed. Instrument should be fine until 
summer shutdown when we will perform a more comprehensive leak test and 
possibly cryotiger servicing.

Agile operated without issues.

Spicam also had a cryotiger related warming. Servicing has been done that 
will hopefully keep it cold until we can replace the coldhead.

TSpec pattern noise, while still present at a low level, is near 
original levels.

Echelle is operational.  Plans for installation of the new grating are 
being developed.  The new ECam had a few growing pains with mask and file 
issues but we believe these are now resolved.

NIC-FPS has been operating in shared-risk observing status.  The image 
corruption issues are still present on occasion but have not been reported 
to be affecting science observations.

GFP software development by Jon Brinkmann et al. is continuing


o spring, summer instrument upgrades/repairs - Klaene, Ketzeback
  (new DIS grating, new echelle grating, Spicam contamination,
   possible NICFPS repair, GFP upgrade status)

See the instrument update page for details:

The new B400 grating was installed in DIS on 3/28/12.  Performance appears 
basically the same, but users should be sure to take new calibrations.

The echelle grating replacement will be scheduled during summer shutdown. 
Users should try to finish programs before end of July if they are 
concerned about possible changes.  We expect better performance with the 
new grating.

Spicam is running warm, probably due to contamination in the cryotiger 
cooling.  We will be taking it out of service for most of the month of May 
to try to fix this problem.  Users should not plan on using Spicam during 
May.  Also, users should take darks during April and be aware that the 
instrument performance may be degraded.

We are awaiting a proposal from UVa to try to address the NICFPS issue of 
returning bad images about 1% of the time.  The repair mission would 
likely take at least 6 months, possibly during Q3 and Q4 (July-Dec) 2012.

===> ACTION: Users committee members to discuss implications of taking 
NICFPS out of service for six months.  Are users happy with the current 
state, is it worth trying to fix?

GFP status is still shared risk, waiting on TUI interface to be


o summary of NIIS visiting instrument run

With significant effort by APO staff, especially Ed Leon and Bill 
Ketzeback, and Matt Nelson (UVa), the instrument was able to be mounted on 
the telescope and data were obtained, albeit with some noise and software 
issues.  Only a few images from APO requests were taken; these will be 
made available to everyone after Mike Pierce (the Wyoming PI) submits a 
written report on the run and provides the data.  It was a successful 
proof of concept for the instrument.


o time trade policies

Suzanne asked that users send time trade requests to her, even if trades 
are within the same institution, and especially if trades are between 
institutions.  Both institutional schedulers need to approve trades.  At 
least 1-2 weeks notice is appreciated. Also, if users wish to give time 
back (e.g. an hour at the end of a half night) please notify Suzanne so 
she can make it available to users who have requested short observations 
and/or to engineering staff.  Schedule updates are frequently made after 
the quarter starts to accomodate trades, users should check the web page. 
An up-to-date schedule is helpful to both the observing specialists and 
the day staff so that telescope activities can be planned efficiently.


o APO future discussion

Suzanne reported that the After Sloan 3 (AS3) project is soliciting seed 
money and expressions of interest in participation.  People should contact 
the AS3 Director, Mike Blanton (NYU).

At the November 2011 ARC Board meeting, a "new partners" committee was 
formed to investigate ways to advertise for new ARC partners in the coming 
years.  This committee has met and is preparing to send out a letter to 
astronomy departments across the US, and some foreign institutions. 
Input including contact names at particular institutions that may be 
interested is welcome, please send information to Suzanne and Rene 
Walterbos (Chair of the ARC Board).  Among the new options (in addition to 
standard ARC membership) will be the possibility of leasing time, and of 
several smaller institutions forming a participation group to purchase a 
block of time together. Current ARC partners who wish to augment their 
time allocation should also contact Suzanne and Rene.

===> ACTION: Send new partner suggestions to Suzanne and Rene (before next 


   o other business -- none.



[open from previous meeting]:

===> ACTION:  Users and Users Committee members to provide Mark and 
Suzanne with their opinions about the impacts of leaving the TripleSpec 
slit wheel permanently set to the 1.1 arcsecond position.

All users committee members reported that leaving the TripleSpec slit 
wheel set to the 1.1" slit is fine.  Recommendation is therefore that the 
instrument not be opened to try to fix this problem.  This action is 

[new from this meeting]:

===> ACTION: Users committee members to discuss implications of taking 
NICFPS out of service for six months.  In particular, are users happy with 
the current state, is it worth trying to fix?

===> ACTION: Send new partner suggestions to Suzanne and Rene (before next 


Next meeting:  The next Users Committee phonecon will be on Monday May 7 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to 
the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1269 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to